# A note on real algebraic maps which are topologically special generic maps

### 09 Jan 2024 v2

@@ Abstract.
This paper also studies real algebraic geometry. Complex algebraic geometry is presenting one of central topics in algebraic geometry. Real algebraic geometry has also another nice history. [24, 35] show classical and important studies and [1, 2, 16, 17] explain fundamental notions and arguments and advanced studies, for example. Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (canonical projections of) spheres.
@@
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is of course a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [23] for example. Hereafter, for a manifold $X$, $\partial X$ denote the boundary of it.
@@
Let $c:X\rightarrow Y$ be a differentiable map between differentiable manifolds $X$ and $Y$. $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$. $c(x)$ is called a singular value of $c$. Let $S(c)$ denote the set of all singular points of $c$. In our paper, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth. A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$. This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$. A canonical projection of the unit sphere $S^{k-1}$ is its restriction.
@@
1.2. Our main results.
Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies. In short two smooth maps are mutually close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies. They respect differentials of degrees at most $r$: $r$ is a non-negative integer of course. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. We omit rigorous exposition. See [8] for systematic exposition. Diffeomorphisms are defined as smooth maps being also homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before of course. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear isomorphisms, defined naturally.
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result.
@@
    We have a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$. By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-N(\partial W_{f})$ whose fiber is the unit sphere.
@@
    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}$.
@@
The next section proves Main Theorems. The third section explains about meanings of our main result and related topics by some remarks.
@@
In [15], followed by [13], the author has redefined notions and consider specific reconstruction and we also abuse these notions for example. Hereafter, for a manifold $X$, ${\rm Int}\ X$ denote the interior $X-\partial X$ of it.
@@
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$. Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$. We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
@@
    The restriction of the interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ to $\bar{N}-N(\partial\bar{N})\subset\bar{N}-{\rm Int}\ N(\partial\bar{N})$ and the bundle over $\bar{N}-N(\partial\bar{N})$ in Proposition 1 are isomorphic and naturally identified.
@@
    We also consider the restriction of bundle ${\bar{M}}_{\rm I}$ in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here. We consider their natural identification. This coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
@@
[15] has first introduced such a class of special generic maps. We first name the class in this way first in [12]. Of course in our present paper, we do not need to understand this notion or [12, 15] well and we do not assume their arguments or results.
@@
    $M_{0}$ is also represented as the zero set of some real polynomial.
@@
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” and have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$ and whose structure group is reduced to a group consisting of linear transformations of $D^{m-n+1}$: in fact a manifold homeomorphic to $D^{k}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined except for $m-n+1=k\neq 4$.
@@
    Under the structure group, the bundle of (1) here is in fact trivial.
@@
      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ homeomorphic to $D^{m-n+2}$. Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
@@
Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered.
@@
We prove our main result. First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
@@
By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$. This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$. Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n}}(\bar{N})$. Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$. We may also regard that ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ by considering a suitable isotopy.
@@
In STEP 1, we have seen that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function. Around a point of the form $(x,y)\in\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way. By considering the partial derivative for some variable $x_{i}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value. We apply the implicit function theorem to see that
@@
is a smooth closed manifold, non-singular real algebraic manifold of dimension $m$ and the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the standard way.
@@
The resulting map is also special generic by the construction and product organized. By considering the isotopy in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2. This completes the proof of Main Theorems 1 and 2.
@@
We first show Main Theorem 3 (1). First the bundle we discuss here is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). We apply an argument used to show Proposition 1. More precisely, Proposition 2.1 of [28] is a corresponding result. It also refers to [22] there. The structure group is regarded as a group consisting of diffeomorphisms preserving the distance between each point and the origin $0\in{\Delta}^{m-n+1}$ where we can define the origin 00 for ${\Delta}^{m-n+1}$ not diffeomorphic to $D^{m-n+1}$. We explain about this. For an element ${\phi}_{{\Delta}^{m-n+1}}$ of such a group, we consider $\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)$ for $x\in{\Delta}^{m-n+1}$ and $0<t\leq 1$ where $x$ and $tx$ are canonically regarded as vectors. Here $||\frac{1}{t}{\phi}_{{\Delta}^{m-n+1}}(tx)||=||x||$. We consider the limit at $t=0$. For a fixed $x$, we have the value of the differential of ${\phi}_{{\Delta}^{m-n+1}}$ at $x$, seen naturally as a tangent vector at the origin of ${\Delta}^{m-n+1}$. Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”. This means that the structure group reduces to the structure group of a linear bundle whose fiber is $D^{m-n+1}$. This completes the proof of Main Theorem 3 (1).
We see that this bundle is trivial for the structure group.
$W_{0}$ is defined as
@@
From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$. By the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$, we can see that the previous bundle is trivial for the structure group. This completes the proof of Main Theorem 3 (2).
@@
In [10, 11, 14], we consider cases where polynomials are given beforehand. In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark and this does not affect main ingredients of [10, 11, 14]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
@@
More precisely, such studies are on Reeb graphs of these functions. The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages. [26] is a pioneering study considering Reeb graphs. Such tools are fundamental and strong tools in studying manifolds by using graphs. In 2006 [29] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph. [18, 20] show explicit answers. [9] is motivated by such studies. There the author has succeeded in construction of nice smooth functions with prescribed preimages first.
[10] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations first. This is also regarded as a study finding a class of real algebraic functions including the canonical projection of a unit sphere. For the canonical projection of the unit sphere, its Reeb graph consists of exactly one edge and exactly two vertices. For this, [3, 31, 32] are also closely related, help the author to do such studies, and motivate us to study further.

### 21 Jan 2024 v3

@@ Abstract.
This paper also studies real algebraic geometry. Complex algebraic geometry is presenting one of central topics in algebraic geometry. Real algebraic geometry has also another nice history. [34, 46] show classical and important studies and [1, 2, 17, 18] present fundamental notions and arguments and advanced studies, for example. Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (canonical projections of) spheres.
@@
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is of course a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [33] for example. Hereafter, for a manifold $X$, let $\partial X$ denote the boundary of it and ${\rm Int}\ X:=X-\partial X$ denote the interior of it.
@@
Let $c:X\rightarrow Y$ be a differentiable map between differentiable manifolds $X$ and $Y$. $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$. $c(x)$ is called a singular value of $c$ for a singular point $x$ of $c$. Let $S(c)$ denote the set of all singular points of $c$. Here, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth. A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$. This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$. A canonical projection of the unit sphere $S^{k-1}$ is its restriction.
@@
1.2. Our main result and the structure of our paper.
Spaces of smooth maps between given two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies. In short two smooth maps are mutually close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies. They respect differentials of degrees at most $r$: $r$ is a non-negative integer of course. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. We omit rigorous exposition. See [9] for systematic exposition. Diffeomorphisms are defined as smooth maps being also homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before of course. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear transformations, defined naturally.
We present a theorem on fundamental structures of special generic maps. Using the notation, we also present our main result as Main Theorem 1. This is proven in the next section with additional results Main Theorems 2 and 3. The third section is on meanings of our main result and related topics.
@@
    We have a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$. By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-{\rm Int}\ N(\partial W_{f})$ whose fiber is the unit sphere.
@@
    $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$.
@@
In [16], followed by [14], the author has redefined notions and consider specific reconstruction and we also abuse these notions for example.
@@
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $\partial\bar{N}\times[0,1]$. Furthermore, points $x\in\partial\bar{N}$ are identified with $(x,0)\in\partial\bar{N}\times\{0\}$ and $\partial N(\partial\bar{N})-\partial\bar{N}$ is regarded as $\partial\bar{N}\times\{1\}$ naturally and the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The last object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$. We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
@@
    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ in Proposition 1 are isomorphic and naturally identified.
@@
    We also consider the restriction of the smooth bundle in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here. We consider their natural identification. This coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
@@
[16] has first introduced such a class of special generic maps. We first name the class in this way first in [13]. Of course in our present paper, we do not need to understand this notion or [13, 16] well. We do not assume their arguments or results.
@@
    $M_{0}\subset{\mathbb{R}}^{m+1}$ is also represented as the zero set of some real polynomial.
@@
    We can abuse the notation in Proposition 1 by replacing ”$f$” by ”$f_{0}$” for example. We have a smooth bundle whose projection is defined as the composition of the restriction of $q_{f_{0}}$ to the preimage of $N(\partial W_{f_{0}})$ with the canonical projection to $\partial W_{f_{0}}$, whose fiber is homeomorphic to $D^{m-n+1}$, the boundary of whose fiber is diffeomorphic to $S^{m-n}$ and whose structure group is reduced to a group isomorphic to the group consisting of all linear transformations of $D^{m-n+1}$: in fact a smooth manifold which is homeomorphic to $D^{k}$ and whose fiber is diffeomorphic to $\partial D^{k}=S^{k-1}$ is known to be diffeomorphic to $D^{k}$ and this is a linear bundle as defined before except for $m-n+1=k\neq 4$.
@@
    Under the presented structure group, the bundle of (1) here is in fact trivial.
@@
      Prepare the product $\bar{N}\times D^{m-n+1}$ and the product $(\partial\bar{N})\times{\Delta}^{m-n+2}$ for a manifold ${\Delta}^{m-n+2}$ which is homeomorphic to $D^{m-n+2}$ and whose boundary is diffeomorphic to $\partial D^{m-n+2}=S^{m-n+1}$. Note that ${\Delta}^{k}$ is diffeomorphic to $D^{k}$ for $k\neq 4$.
@@
Remember that ”${\Delta}^{k}$ not being diffeomorphic to $D^{k}$” is still undiscovered. If this holds for some ${\Delta}^{k}$, then of course $k$ must be $4$.
@@
First we prove Main Theorems 1 and 2. STEPs 1–3 are main ingredients.
@@
By considering a small isotopy considered in the class $C^{1}$ starting from the original embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$, we have another smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n},0}:\bar{N}\rightarrow{\mathbb{R}}^{n}$. We can also consider the isotopy in the class $C^{a}$ where $a$ is an arbitrary integer greater than $1$ according to [19]. For this approximation, see also [3, 7].
This fixes the manifold except a small collar neighborhood $N(\partial\bar{N})$ of the boundary $\partial\bar{N}$. Let ${\bar{N}}_{0}\subset{\mathbb{R}}^{n}$ denote the resulting image ${\bar{f}}_{{\mathbb{R}}^{n},0}(\bar{N})$. Let ${\partial}_{j}{\bar{N}}_{0}$ denote all connected components of the boundary $\partial{\bar{N}}_{0}$, indexed by all positive integers $j=1,\cdots,l\geq 1$ smaller than or equal to $l$. We may also regard that ${\partial}_{j}{\bar{N}}_{0}$ is a non-singular real algebraic manifold of dimension $n-1$ and the zero set of a real polynomial $f_{{\partial}_{j}}$ (by considering a suitable small isotopy).
@@
The last part of STEP 1 implies that this is also represented as $\{(x,y)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}\subset{\mathbb{R}}^{m+1}\mid{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}=0\}$.
Around a point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, it is represented as the graph of a canonically obtained smooth function where some $y_{j}$ is regarded as the variable of the target of the function. Around a point of the form $(x,y)\in\partial{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ in this subset, we consider the implicit function theorem for the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$ in a different way.
By considering the partial derivative for some variable $x_{j_{0}}$ in $x=(x_{1},\cdots x_{n})$, we have a positive value. We explain about this more precisely. If $x\in{\partial}_{i_{0}}{\bar{N}}_{0}$ holds, then for some variable $x_{j_{0}}$, we have the value represented as the product of the partial derivative of $f_{{\partial}_{i_{0}}}(x)$ (for $x_{j_{0}}$) and the product of the values of the remaining polynomials $f_{{\partial}_{j}}(x)$. This is thanks to elementary formula in calculus and the property of the zero sets of our polynomials. Remember that $f_{{\partial}_{j}}(x)$ is not zero for any $j$ except $j=i_{0}$. The first value is not zero since each ${\partial}_{i}{\bar{N}}_{0}$ is obtained by some suitable approximation of an original connected component being also a smooth compact manifold with no boundary: remember that this is also obtained via isotopy in the $C^{1}$ class. The second value is not zero by the property of the zero sets of the polynomials.
We apply the implicit function theorem to see that
@@
is a smooth, closed and connected manifold and non-singular real algebraic manifold of dimension $m$. Remember that this is also the zero set of the polynomial ${\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))-{\Sigma}_{j=1}^{m-n+1}{y_{j}}^{2}$. The map $f_{0}$ is obtained as the composition of the canonical embedding of $M_{0}\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$.
@@
We can map each point of the form $(x,y)\in{\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,\frac{g(t)}{||y||}y)$ according to the following rule: $y$ and $\frac{g(t)}{||y||}y$ are regarded as vectors in the canonical way.
@@
The resulting map is also special generic by the construction. It is also product-organized. By considering the isotopy (in the $C^{1}$ class) in STEP 1, we have a special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ with the smooth real algebraic map $f_{0}$ enjoying all desired properties in Main Theorems 1 and 2.
This completes the proof of Main Theorems 1 and 2.
@@
We define $W_{0}$ as
@@
From this we can see that $W_{0}-S(f_{0})$ is regarded as the total space of a trivial smooth bundle over ${\rm Int}\ {\bar{N}}_{0}$ whose fiber is diffeomorphic to $D^{m-n+1}$.
We investigate the bundle in Main Theorem 3 (1) and (2). Let the bundle denoted by $M_{{\rm Boundary},0}:=M_{{\rm B},0}$. First the bundle $M_{{\rm B},0}$ is a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). We apply an argument used to show Proposition 1 in a revised style. More precisely, Proposition 2.1 of [39] is a corresponding result. It also refers to [32] there. We consider the restriction of the bundle $W_{0}-S(f_{0})$ over ${\rm Int}\ {\bar{N}}_{0}$ to ${\bar{N}}_{0}-N(\partial{\bar{N}}_{0})\subset{\rm Int}\ {\bar{N}}_{0}$. We remove the total space of this bundle obtained by the restriction from $W_{0}$. We have a smooth bundle $W_{1}$ over $\partial W_{f_{0}}$ whose fiber is homeomorphic to $D^{m-n+2}$. This is diffeomorphic to $D^{m-n+2}$ in the case $m-n+2\neq 4,5$. This is also thanks to Ehresmann’s fibration theorem (or a relative version of this). Hereafter we apply this implicitly in several scenes. The bundle $M_{{\rm Boundary},0}=M_{{\rm B},0}$ is regarded as a subbundle of the bundle $W_{1}$ over $\partial W_{f_{0}}$. The structure group of the bundle $M_{{\rm B},0}$ is restricted to the structure group of the (bundle canonically regarded as a) subbundle defined on the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ for each $0<t\leq 1$. For each $0<t\leq 1$, it is a subbundle of the bundle obtained naturally as the subbundle of $W_{1}$ by restricting the total space to the preimage of $\partial{\bar{N}}_{0}\times\{t\}$ similarly. For each two obtained smooth bundles, fibers are regarded as $\partial D^{m-n+1}$ and $D^{m-n+1}$, respectively. The former bundle is also obtained by restricting the fiber to the boundary $\partial D^{m-n+1}$ of course. We consider an element ${\psi}_{t}$ of the structure group of the subbundle $M_{{\rm B},0,t}$ of $M_{{\rm B},0}$ defined for each $0<t\leq 1$ before. This element can be regarded as an element extended to an element preserving the distance between each point and the origin $0\in D^{m-n+1}$. This is also regarded as an element of the structure group of the subbundle $W_{1,t}$ of $W_{1}$ defined for each $0<t\leq 1$ before. For such an element, we can consider $\frac{1}{s}{\psi}_{t}(sv)$ for $v\in D^{m-n+1}$ which is not the origin and $0<s\leq 1$. Of course $v$, $sv$ and $\frac{1}{s}{\psi}_{t}(sv)$ are canonically regarded as vectors and hereafter we do not explain about this rule explicitly. Here $||\frac{1}{s}{\psi}_{t}(sv)||=||v||$. We consider the limit at $s=0$. For a fixed $v$, we have the value of the differential of ${\psi}_{t}$ at $v$, seen naturally as a tangent vector at the origin of ${D}^{m-n+1}$. Thus we naturally have an element regarded as a linear transformation of ”the unit disk $D^{m-n+1}$”. This means that the structure group reduces to the group isomorphic to the structure group of a linear bundle whose fiber is $D^{m-n+1}$. These arguments and the structure of the map around $S(f_{0})$ and $\partial{\bar{N}}_{0}$ complete the proof of Main Theorem 3 (1). We can also see that the bundle $M_{{\rm B},0}$ is trivial for the structure group. This completes the proof of Main Theorem 3 (2).
@@
In [11, 12, 15], we consider cases where polynomials are given beforehand. In short, only STEP 2 is considered and related short exposition on STEP 3 is presented where the latter presentation is a kind of short remark. This does not affect main ingredients of [11, 12, 15]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
@@
As another problem which is in some sense similar to this problem, it is also very difficult to know whether two homeomorphisms $\Phi$ and $\phi$ in Main Theorem 1 can be chosen as diffeomorphisms in general.
@@
More precisely, such studies are on Reeb graphs of these functions. The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages. The vertex set of the Reeb graph consists of all elements representing connected components containing some singular points of the function.
[36] is a pioneering study considering Reeb graphs. Such tools are fundamental and strong tools in studying manifolds by using graphs. In 2006, [40] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph. [20, 30] show explicit answers. [10] is motivated by such studies. There the author has succeeded in construction of nice smooth functions with prescribed preimages first.
[11] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations first. This is also regarded as a study finding a class of real algebraic functions including the canonical projection of a unit sphere. For the canonical projection of the unit sphere, its Reeb graph consists of exactly one edge and exactly two vertices. For this, [3, 42, 43] are also closely related, help the author to do such studies, and motivate us to study further. They present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ and regions in ${\mathbb{R}}^{2}$ surrounded by non-singular real algebraic curves.
Related to Remark 3, we add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [22, 23, 24, 25, 26, 27, 28, 29]. See also [37]. This is also closely related to [3]. This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory and this mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs. We do not explain about elementary notions on graphs precisely. We except related knowledge. We also note that we omit several rigorous discussions in Remark 4 and that for our main ingredients of the present paper we do not need to understand this rigorously.
Remark 4.
[3] considers a finite graph $K$ enjoying the following properties.
  • We can embed $K$ by a piecewise smooth embedding $e:K\rightarrow{\mathbb{R}}^{2}$.
  • The restriction of the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to each edge is an smooth embedding.
  • The degree of each vertex is $1$ or $3$.
  • At the vertex where the function ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ has a local extremum, the degree of the vertex is $1$.
An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves. The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of connected components of all preimages of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$ under the natural equivalence relation. Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\overline{R}}$. The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ containing $e(K)$ and its Poincaré-Reeb graph is isomorphic to and naturally regarded as $e(K)$. Moreover, we can have the situation in the following.
  1. (1)
    The restriction of ${\pi}_{2,1}\circ e:K\rightarrow\mathbb{R}$ to the boundary $\partial\overline{R_{K}}$ is a Morse function such that at distinct singular points the values are always distinct.
  2. (2)
    $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial. This is also obtained by approximation of the Weierstrass type in the class $C^{2}$.
According to arguments similar to one in ”Special Case 5” of [17], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.